Find dy/dx and d²y/dx² of;


x = theta + cos theta, y = 1 + sin theta


; theta = pi/6

1
Expert's answer
2022-01-26T18:24:51-0500

i will use parameter t instead of theta

for parametric function

first derivative "y'_x" can be found as "{\frac {y'_t} {x'_t}}"

second derivative "y''_{xx}" can be found as "{\frac {(y'_x)'_t} {x'_t}}"


"y'_t=cos(t)"

"x'_t=1-sin(t)"

so,

"y'_x={\frac {cos(t)} {1-sin(t)}}\implies y'_x({\frac {\pi} 6})={\frac {cos({\frac {\pi} 6})} {1-sin({\frac {\pi} 6})}}=\sqrt3"


"(y'_x)'_t={\frac {-sin(t)+sin^2(t)+cos^2(t)} {(1-sin(t))^2}}={\frac 1 {(1-sin(t))}}"

so,

"y''_{xx}={\frac 1 {(1-sin(t))^2}}\implies y''_{xx}({\frac {\pi} 6})={\frac1 {(1-sin({\frac {\pi} 6}))^2}}=4"